How long (in years and months) will it take for an investment to double at 4% compounded monthly?
Answer by Alan3354(69443)

How long (in years and months) will it take for an investment to double at 4% compounded monthly?
----------------
If it's 4% per year,
2 = (1 + 0.04/12)^t, t in months
log(2) = t*log(1+0.04/12)
t = log(1+0.04/12)/log(2)
Use a calculator from there